### v1

@@ 5 Discussion
In this paper, we have developed a systematic Monte Carlo approach to randomly sample $d$-dimensional lattice polytopes from a fixed set of points. With the properly derived weight factors, one can study the statistics of these lattice polytopes. We applied this method to study toric 3d bases for 4d F-theory models. Such statistical methods can also potentially be applied to other combinatorics problems. For instance, we observe that the number of polytopes satisfying some general classes of conditions obey a Gaussian distribution with respect to the number of vertices $m$. One can, for instance, see this phenomenon in the Kreuzer-Skarke set of reflexive polytopes.
@@
For $d=3$, we have first identified a set of “maximal” 3d boxes in which one can sample the base polytopes. These boxes give rise to a set of minimal dual $\mathcal{G}$-polytopes, which can also be used as starting points of blow-ups in 3d toric minimal model program. While we did not construct the complete set of maximal 3d boxes and corresponding minimal 3d $\mathcal{G}$-polytopes, we believe that we have identified the majority of these boxes, and the results from the boxes we analyzed should form a representative sample; we leave a more complete analysis of the set of 3d boxes as a challenge for future work. Sampling the polytopes in the maximal 3d boxes gives a distribution on the set of toric threefold bases for 4d F-theory models, in which only polytope data is maintained, so that bases that are birationally equivalent are identified. This groups together bases with similar physics and simplifies the combinatorial structure of this part of the F-theory landscape, giving somewhere in the range of $10^{85}$$10^{90}$ distinct base polytopes associated with 3d toric varieties that can support elliptic Calabi-Yau fourfolds. On the other hand, this ensemble also broadens the class of toric bases somewhat beyond those considered previously. In particular, note that that the base polytopes we found through this approach typically lead to bases with terminal singularities. The physical impact of these base singularities in 4d F-theory should be the subject of future investigations. The base polytopes also possess non-minimal codimension-two (4,6) loci in the elliptic fibrations over these bases, which lead to strongly coupled F-theory sectors at least at a high energy scale Apruzzi:2018oge ; Tian:2018icz . It would also interesting to investigate the phenomenological impact of such sectors.
@@
In this new measure for sampling 3d bases, we also found that the presence of $E_{6}$ or $E_{7}$ gauge groups that can be useful in Standard Model building Li:2021eyn ; Li:2022aek ; Li:2023dya are not that uncommon, in particular if we are away from the maximal box $\mathcal{B}_{6}(84,516)$; see Figure 22 and 23. For such Standard Model constructions, however, typically a non-toric base is necessary, so for explicit model building one either needs an alternative way to realize hypercharge flux breaking and the Higgs sectors in order to actually construct “generic” MSSM on these toric base geometries, or to extend this analysis to non-toric bases. More generally, however, assuming that the distribution of gauge groups is not too different for more general (non-toric) base geometries, these statistics support the hypothesis that rigid $E_{6}$ and $E_{7}$ gauge groups are reasonably common in the 4D F-theory landscape, at least as features of birational equivalence classes of bases.
@@
As in previous studies, we find that of the rigid/non-Higgsable gauge factors that contain the standard model, $E_{8}$ is by far the most prevalent, even without accounting for flux multiplicities. Thus, despite the challenges posed by the presence of strongly coupled matter at SCFT points Kim:2017toz ; Apruzzi:2018oge ; Tian:2018icz , rigid $E_{8}$ factors remain the most promising and abundant place to find Standard Model-like physics in the landscape of 4D F-theory models.
@@
While the set of non-toric threefold bases is much less well understood than the set of toric bases, the approach of this paper, which groups together birationally equivalent threefold bases, may be helpful as well in the non-toric context.
@@ Acknowledgements.
We would like to thank Shing Yan Li and Richard Nally for helpful discussions. YNW and YHY are supported by National Natural Science Foundation of China under Grant No. 12175004, No. 12422503 and by Young Elite Scientists Sponsorship Program by CAST (2023QNRC001, 2024QNRC001). The work of WT was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567. The work is supported by High-performance Computing Platform of Peking University. This work also made use of resources provided by subMIT at the MIT physics Department bendavid2025submitphysicsanalysisfacility .

### v2

@@ 5 Discussion
In this paper, we have developed a systematic Monte Carlo approach to randomly sample $d$-dimensional lattice polytopes from a fixed set of points. With the properly derived weight factors, one can study the statistics of these lattice polytopes. We applied this method to study toric 3d bases for 4d F-theory models. Such statistical methods can also potentially be applied to other combinatorics problems. For instance, we observe that the number of polytopes satisfying some general classes of conditions obey a Gaussian distribution with respect to the number of vertices $m$. One can, for instance, see this phenomenon in the Kreuzer-Skarke set of 4d reflexive polytopes.
A future direction would be applying the Monte Carlo methods in this paper to generate reflexive polytopes of higher dimensions. For instance, one can try to take the $k=1$ boxes and sample $n$ points in them. If the resulting polytope is not reflexive, one takes the $k=1$ polar dual and only keeps the lattice points in that. The resulting lattice polytope would be a reflexive polytope by definition, and all reflexive $d$-polytopes can be sampled in this form. Additional filtering and redundancy factors should be included for such a statistical analysis of reflexive polytopes to work. Note, however, that the full $k=1$ set of polytopes, which includes Calabi-Yau hypersurface threefolds with various singularity types, may also be of direct physical interest AGNT.
@@
For $d=3$, we have first identified a set of “maximal” 3d boxes in which one can sample the base polytopes. These boxes give rise to a set of minimal dual $\mathcal{G}$-polytopes, which can also be used as starting points of blow-ups in 3d toric minimal model program. While we did not construct the complete set of maximal 3d boxes and corresponding minimal 3d $\mathcal{G}$-polytopes, we believe that we have identified the majority of these boxes, and the results from the boxes we analyzed should form a representative sample; we leave a more complete analysis of the set of 3d boxes as a challenge for future work — this may be possible following an approach similar to that used in Avram:1997rs for the $k=1$ case of maximal boxes. Sampling the polytopes in the maximal 3d boxes gives a distribution on the set of toric threefold bases for 4d F-theory models, in which only polytope data is maintained, so that many bases that are birationally equivalent (through, e.g., flips and flops), with the same divisor and gauge group structures are identified. This groups together bases with similar physics and simplifies the combinatorial structure of this part of the F-theory landscape, giving somewhere in the range of $10^{85}$$10^{90}$ distinct base polytopes associated with 3d toric varieties that can support elliptic Calabi-Yau fourfolds. On the other hand, this ensemble also broadens the class of toric bases somewhat beyond those considered previously. In particular, note that that the base polytopes we found through this approach typically lead to bases with terminal singularities on the base, such as the $U(3)$ orbifold singularity $\mathbb{C}^{3}/\mathbb{Z}_{2}(1,1,1)$ Morrison:2004fr. The physical impact of these base singularities in 4d F-theory should be the subject of future investigations. The base polytopes also possess non-minimal codimension-two (4,6) loci in the elliptic fibrations over these bases, which lead to strongly coupled F-theory sectors at least at a high energy scale Apruzzi:2018oge; Tian:2018icz. It would also be interesting to investigate the phenomenological impact of such sectors.
As in previous studies of the F-theory landscape, one central takeaway from this analysis is the ubiquitous presence of many rigid gauge group factors in virtually all of the geometries sampled. In addition to providing natural gauge sectors for realizing the Standard Model, most models have many additional “hidden sector” SU(2), $G_{2},F_{4}$, and $E_{8}$ gauge factors that could act as dark matter. Furthermore, the strongly coupled (4, 6) sectors that arise in the geometries considered here should also be considered as potential dark matter candidates. [^11] We would like to thank Jean Du Plessis for discussion on this point.
@@
In this new measure for sampling 3d bases, we found in particular that the presence of $E_{6}$ or $E_{7}$ gauge groups that can be useful in Standard Model building Li:2021eyn; Li:2022aek; Li:2023dya are not that uncommon, in particular if we are away from the maximal box $\mathcal{B}_{6}(84,516)$; see Figure 22 and 23. For such Standard Model constructions, however, typically a non-toric base is necessary, so for explicit model building one either needs an alternative way to realize hypercharge flux breaking and the Higgs sectors in order to actually construct “generic” MSSM on these toric base geometries, or to extend this analysis to non-toric bases. More generally, however, assuming that the distribution of gauge groups is not too different for more general (non-toric) base geometries, these statistics support the hypothesis that rigid $E_{6}$ and $E_{7}$ gauge groups are reasonably common in the 4d F-theory landscape, at least as features of polytope equivalence classes of bases.
@@
As in previous studies, we find that of the rigid/non-Higgsable gauge factors that contain the Standard Model, $E_{8}$ is by far the most prevalent, even without accounting for flux multiplicities. Thus, despite the challenges posed by the presence of strongly coupled matter at SCFT points Kim:2017toz; Apruzzi:2018oge; Tian:2018icz, rigid $E_{8}$ factors remain the most promising and abundant place to potentially find Standard Model-like physics in the landscape of 4d F-theory models.
@@
While the set of non-toric threefold bases is much less well understood than the set of toric bases, the approach of this paper, which groups together threefold bases related by flops and flips, may be helpful as well in the non-toric context.
@@ Acknowledgements.
We would like to thank Jean Du Plessis, Shing Yan Li and Richard Nally for helpful discussions. YNW and YHY are supported by National Natural Science Foundation of China under Grant No. 12175004, No. 12422503 and by Young Elite Scientists Sponsorship Program by CAST (2023QNRC001, 2024QNRC001). The work of WT was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567. The work is supported by High-performance Computing Platform of Peking University. WT would like to thank the Santa Fe Institute, where some of this work was completed. This work also made use of resources provided by subMIT at the MIT physics Department bendavid2025submitphysicsanalysisfacility.